# Beyond Rate Coding: Surrogate Gradients Enable Spike Timing Learning in Spiking Neural Networks

### v2

@@ 5 Discussion
Surrogate gradient descent and related methods (Neftci et al., 2019; Shrestha and Orchard, 2018) have proven themselves exceptionally capable at the challenging task of training non-differentiable spiking neural networks to carry out challenging tasks. However, the question remains whether or not the solutions found using these algorithms cover the full range of capabilities of spiking neural networks, or whether they are biased to a particular subfamily of solutions. In both cases, are the solutions found similar to what biological neural networks would learn? We addressed the particular question of the extent to which timing rather than rate of action potentials can be used.
In principle, a timing-based code can be vastly more efficient than a rate-based one. Suppose spike times were accurate to within 1 ms and there was no restriction on firing rate, then for a $k$ ms window, spike count can encode $k+1$ different values whereas with spike timing we can encode $2^{k}$ values. In practice, the presence of neural noise and the difficulty of decoding the full range of possible timing codes will limit this. Some of the aspects of spike timing that certainly are robustly decodable include: inter-spike intervals (the time between spikes), cross-neuron coincidences, and at least some more complex spatio-temporal spike patterns. In our results, we have shown that surrogate gradient descent-based methods are able to extract information using all of these aspects of timing-based codes, in both controlled synthetic datasets and messier datasets based on recorded speech. In addition, under various forms of spike time perturbation they degrade gracefully in the majority of cases.
@@
Finally, we address the question of whether or not these algorithms find solutions that are biologically feasible by considering what happens when you train a network on a set of sounds, and then test it by reversing those sounds. Since humans perform badly at this task, we would expect models that learn in a way that is more similar to humans to also perform badly. We find that spiking neural networks both with and without delays are negatively impacted by time reversal, but that networks with delays are much more significantly impacted. This suggests that networks with delays may be more biologically relevant, at least in the case of modeling the auditory system, where we know that timing information is important.
In future work, it would be interesting to investigate other aspects of spike timing-based codes, for example temporal order (A fires before B), or a more fine-grained analysis of which spatio-temporal spike patterns can be learned. It would also be interesting to devise further datasets of real-world complexity in which temporal information is substantially more important than rate-based information.
@@
The code for loading the variant SHD and SSC datasets is available on the GitHub repository: https://github.com/neural-reckoning/temporal-shd. The datasets are archived on Zenodo(Yu et al., 2025) (https://doi.org/10.5281/zenodo.16153275).

### v3

@@ 5 Discussion
Surrogate gradient descent and related methods (neftci2019surrogate; shrestha2018slayer) have proven themselves exceptionally capable at the challenging task of training non-differentiable spiking neural networks to carry out challenging tasks.
However, the question remains whether or not the solutions found using these algorithms cover the full range of capabilities of spiking neural networks, or whether they are biased to a particular subfamily of solutions. In both cases, are the solutions found similar to what biological neural networks would learn? We addressed the particular question of the extent to which timing rather than rate of action potentials can be used.
In principle, a timing-based code can be vastly more efficient than a rate-based one. This comparison assumes a discretized time representation with finite temporal precision, as used in both neuromorphic datasets and most biological modelling contexts. Suppose spike times were accurate to within 1 ms and there was no restriction on firing rate, then for a $k$ ms window, spike count can encode $k+1$ different values whereas with spike timing we can encode $2^{k}$ values. We emphasize that this comparison is purely illustrative under discretized time assumptions and is not meant to reflect achievable operating regimes in practice. In practice, the presence of neural noise and the difficulty of decoding the full range of possible timing codes will limit this. Some of the aspects of spike timing that certainly are robustly decodable include: inter-spike intervals (the time between spikes), cross-neuron coincidences, and at least some more complex spatio-temporal spike patterns. In our results, we have shown that surrogate gradient descent-based methods are able to extract information using all of these aspects of timing-based codes, in both controlled synthetic datasets and messier datasets based on recorded speech. In addition, under various forms of spike time perturbation they degrade gracefully in the majority of cases.
@@
Finally, we address the question of whether or not these algorithms find solutions that are biologically feasible by considering what happens when you train a network on a set of sounds, and then test it by time-reversing the spike trains. Since humans perform badly on time-reversed sound, we would expect models that learn in a way that is more similar to humans to also perform badly. We find that spiking neural networks both with and without delays are negatively impacted by time reversal, but that networks with delays are much more significantly impacted.
In future work, it would be interesting to investigate other aspects of spike timing-based codes, for example temporal order (A fires before B), or a more fine-grained analysis of which spatio-temporal spike patterns can be learned. It would also be interesting to devise further datasets of real-world complexity in which temporal information is substantially more important than rate-based information. Finally, it would also be interesting to understand how and when such networks can generate explicit spike-timing codes at the output.
@@
All code in this work is available on the GitHub repository: https://github.com/neural-reckoning/temporal-shd. The datasets are archived on Zenodo(yu2025temporalshd) (https://doi.org/10.5281/zenodo.16153275).